OpenStudy (anonymous):

Replace ? with =, >, or < to make the statement true. 16 ÷ 4 + 12 ? 4 + 24 ÷ 2

OpenStudy (anonymous):

=

OpenStudy (anonymous):

First answer to : 16/4=?

OpenStudy (anonymous):

so we have : 4 and have : 4+12 ? 4+24/2 24/2=12 and have : 4+12 ? 4+12 so ? is =

OpenStudy (anonymous):

A. x = 1, y = 1 B. x = 1, y = 2 C. x = 2, y = 1 D. x = 4, y = 3

OpenStudy (anonymous):

and where is the question ???